# Features of the Partition Function of a $\Lambda>0$ Universe

### v1

@@ B.2 Linearised graviton
~$\tilde{\mathcal{I}}_{11}=\frac{4}{\varepsilon^{4}}+\frac{4}{\varepsilon^{2}}-12\log A+\frac{599}{120}\log\frac{e^{\gamma}\varepsilon}{2}+\zeta^{\prime}(-3)+-\frac{13}{3}+\frac{\log 3}{2}+8\log 2~{}.$ (B.4)
@@
~$\tilde{\mathcal{I}}_{12}=\frac{8}{\varepsilon^{4}}+\frac{2}{\varepsilon^{2}}+\frac{13}{48}+\frac{7}{30}\log\frac{\varepsilon e^{\gamma}}{2}+\frac{1}{2}\sum_{\pm}P(\hat{\delta}-\Delta_{\pm})\zeta^{\prime}(0,\Delta_{\pm})~{},$ (B.5)
@@
~$\tilde{\mathcal{I}}_{22}=\frac{8}{\varepsilon^{4}}+\frac{2}{\varepsilon^{2}}+\frac{13}{48}+\frac{607}{30}\log\frac{e^{\gamma}\varepsilon}{2}+\frac{1}{2}\sum_{\pm}P(\hat{\delta}-\Delta_{\pm})\zeta^{\prime}(0,\Delta_{\pm})~{},$ (B.6)
@@
~$\tilde{\mathcal{I}}_{22}-\tilde{\mathcal{I}}_{21}=20\log\frac{\sqrt{3}e^{\gamma}\varepsilon}{2}=\frac{1}{2}\times 20\log\frac{2\Lambda}{\Lambda_{\text{u.v.}}}~{}.$ (B.7)
@@
~$\tilde{\mathcal{I}}_{23}=\frac{8}{\varepsilon^{4}}-\frac{16}{\varepsilon^{2}}-\frac{1741}{60}\log e^{\gamma}\varepsilon+12\log A+2\zeta^{\prime}(-3)+\frac{22}{3}+\frac{1261}{60}\log 2-10\log 3~{}.$ (B.8)
@@ C.2 Tensor spectrum
with $n_{1},n_{2}\in\mathbb{Z}^{+}+1$. In addition there is a single negative tensor eigenmode with eigenvalue $-2\Lambda$, and nine tensor eigenmodes with eigenvalue $2\Lambda$. The negative mode corrresponds to changing the relative size of the two-spheres without breaking any spherical symmetry.
@@ C.3 One-loop contribution
For the case at hand, $\mu^{2}\Lambda^{-1}=-1$, such that the $\nu_{n_{1}}$ are irrational. Unlike what happens for $S^{4}$ and $\mathbb{C}P^{2}$, the finite piece cannot be reduced to a simple expression in terms of derivatives of the Riemann zeta function. On the other hand, since we have an infinite sum over $SO(2,1)$ characters in (C.7) furnishing the principal series irreducible representation with $\Delta_{n_{1}}=\tfrac{1}{2}+i\nu_{n_{1}}$, perhaps we can view these as enhanced (possibly non-unitary) contributions to the edge mode sector of (1.4), itself built from $SO(2,1)$ characters (see also lawlochab ).

### v3

@@ B.2 Linearised graviton
$\tilde{\mathcal{I}}_{11}=\frac{4}{\varepsilon^{4}}+\frac{4}{\varepsilon^{2}}-12\log A+\frac{599}{120}\log\frac{e^{\gamma}\varepsilon}{2}+\zeta^{\prime}(-3)+-\frac{13}{3}+\frac{\log 3}{2}+8\log 2~.$ (B.4)
@@
$\tilde{\mathcal{I}}_{12}=\frac{8}{\varepsilon^{4}}+\frac{2}{\varepsilon^{2}}+\frac{13}{48}+\frac{7}{30}\log\frac{\varepsilon e^{\gamma}}{2}+\frac{1}{2}\sum_{\pm}P(\hat{\delta}-\Delta_{\pm})\zeta^{\prime}(0,\Delta_{\pm})~,$ (B.5)
@@
$\tilde{\mathcal{I}}_{22}=\frac{8}{\varepsilon^{4}}+\frac{2}{\varepsilon^{2}}+\frac{13}{48}+\frac{607}{30}\log\frac{e^{\gamma}\varepsilon}{2}+\frac{1}{2}\sum_{\pm}P(\hat{\delta}-\Delta_{\pm})\zeta^{\prime}(0,\Delta_{\pm})~,$ (B.6)
@@
$\tilde{\mathcal{I}}_{22}-\tilde{\mathcal{I}}_{21}=20\log\frac{\sqrt{3}e^{\gamma}\varepsilon}{2}=\frac{1}{2}\times 20\log\frac{2\Lambda}{\Lambda_{\text{u.v.}}}~.$ (B.7)
@@
$\tilde{\mathcal{I}}_{23}=\frac{8}{\varepsilon^{4}}-\frac{16}{\varepsilon^{2}}-\frac{1741}{60}\log e^{\gamma}\varepsilon+12\log A+2\zeta^{\prime}(-3)+\frac{22}{3}+\frac{1261}{60}\log 2-10\log 3~.$ (B.8)
@@ C.2 Tensor spectrum
with $n_{1},n_{2}\in\mathbb{Z}^{+}+1$. In addition there is a single negative tensor eigenmode with eigenvalue $2\Lambda$, and nine tensor eigenmodes with eigenvalue $2\Lambda$. The negative mode corrresponds to changing the relative size of the two-spheres without breaking any spherical symmetry.
@@ C.3 One-loop contribution
For the case at hand, $\mu^{2}\Lambda^{-1}=-1$, such that the $\nu_{n_{1}}$ are irrational. Unlike what happens for $S^{4}$ and $\mathbb{C}P^{2}$, the finite piece cannot be reduced to a simple expression in terms of derivatives of the Riemann zeta function. On the other hand, since we have an infinite sum over $SO(2,1)$ characters in (C.7) furnishing the principal series irreducible representation with $\Delta_{n_{1}}=\tfrac{1}{2}+i\nu_{n_{1}}$, perhaps we can view these as enhanced (possibly non-unitary) contributions to the edge mode sector of (1.4), itself built from $SO(2,1)$ characters (see also Law:2025yec ; Mukherjee:2025xlt ).